OpenStudy (anonymous):

For f(x) = 4x+1 and g(x)= x^2 - 5 find (g/f)(X)

OpenStudy (anonymous):

Just divide them simply First put the given value of g(x) in the numinator and then put f(x) value in denominator and show me your work.

OpenStudy (anonymous):

i don't know how to do this problem at all can you show me step by step plz

OpenStudy (anonymous):

g/f(x)=g(x)/f(x) Now put the equation of them in it.

OpenStudy (anonymous):

x^2-5/4x+1

OpenStudy (anonymous):

(g/f)(x)=x^2-5/4x+1 U done well... Now look on the equation is can it divide?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

So this is your answer.

OpenStudy (anonymous):

oh nevermind lol.... thank you so much :)

OpenStudy (anonymous):

R u 18years old?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Okay I m also.

OpenStudy (anonymous):

awesome!!!